# Nearly Tight Black-Box Auditing of Differentially Private Machine Learning

### v1

@@ Abstract
This paper presents a nearly tight audit of the Differentially Private Stochastic Gradient Descent (DP-SGD) algorithm in the black-box model. Our auditing procedure empirically estimates the privacy leakage from DP-SGD using membership inference attacks; unlike prior work, the estimates are appreciably close to the theoretical DP bounds. The main intuition is to craft worst-case initial model parameters, as DP-SGD’s privacy analysis is agnostic to the choice of the initial model parameters. For models trained with theoretical $\varepsilon=10.0$ on MNIST and CIFAR-10, our auditing procedure yields empirical estimates of $7.21$ and $6.95$, respectively, on 1,000-record samples and $6.48$ and $4.96$ on the full datasets. By contrast, previous work achieved tight audits only in stronger (i.e., less realistic) white-box models that allow the adversary to access the model’s inner parameters and insert arbitrary gradients. Our auditing procedure can be used to detect bugs and DP violations more easily and offers valuable insight into how the privacy analysis of DP-SGD can be further improved.
@@ 1 Introduction
Differentially Private Stochastic Gradient Descent (DP-SGD) [1] allows training machine learning models while providing formal Differential Privacy (DP) [14] guarantees. DP-SGD is increasingly deployed in production systems [2, 4, 25] and supported in open-source libraries like Opacus [36], TensorFlow [17], and JAX [5]. Progress in privacy accounting and careful hyper-parameter tuning has helped reduce the gap in model performance compared to non-private training [10]. Furthermore, pre-training models on large public datasets and then fine-tuning them on private datasets using DP-SGD can also be used to improve utility [10].
@@
Implementing DP-SGD correctly can be challenging – as with most DP algorithms [6, 23] – and bugs that lead to DP violations have been found in several DP-SGD implementations [8, 20]. These violations do not only break the formal DP guarantees but can also result in realistic privacy threats [15, 31]. This prompts the need to audit DP-SGD implementations to verify whether the theoretical DP guarantees hold in practice. The auditing procedure usually involves running membership inference attacks [30], whereby an adversary infers whether or not a sample was used to train the model. The adversary’s success estimates the empirical privacy leakage from DP-SGD, and that is compared to the theoretical DP bound. If the former is “far” from the latter, it means that the auditing procedure is loose and may not be exploiting the maximum possible privacy leakage. Whereas, when empirical and theoretical estimates are close, auditing is tight and can be used to effectively identify bugs [26].
@@
Although techniques to tightly audit DP-SGD exist in literature [26, 27], they do so only in active white-box threat models, where the adversary can observe and insert arbitrary gradients into the intermediate DP-SGD steps. However, real-world adversaries cannot always access the model’s inner parameters, let alone insert arbitrary gradients. Furthermore, white-box audits may be hard to perform in practice and at scale, as the algorithm under audit has to be substantially modified, which may not always be possible. For instance, libraries do not always expose intermediate gradients of each sample, e.g., for efficiency reasons. On the other hand, prior work considering the more realistic black-box threat model, where an adversary can only observe the final trained model, has only achieved loose audits [18, 26, 27]. Therefore, it has so far remained an open research question [26] whether or not it is possible to tightly audit DP-SGD in the black-box model.
@@
In this paper, we introduce a novel auditing procedure that achieves tight black-box audits of DP-SGD, even for natural (i.e., not adversarially crafted) datasets. The main intuition behind our auditing procedure is to craft worst-case initial model parameters, as DP-SGD’s privacy analysis is agnostic to the choice of initial model parameters. We show that the gap between worst-case empirical and theoretical privacy leakage in the black-box threat model is much smaller than previously thought [18, 26, 27]. When auditing a convolutional neural network trained on MNIST and CIFAR-10 at theoretical $\varepsilon=10$, our audits achieve an empirical privacy leakage estimate of $\varepsilon_{emp}=6.48$ and $4.96$, respectively, compared to $\varepsilon_{emp}=3.41$ and $0.69$ when the models were initialized to average-case initial model parameters, as done in prior work [10, 18, 26].
@@
In the process, we identify and study the key factors affecting the tightness of black-box auditing, namely: 1) dataset size and 2) gradient clipping norm. Specifically, we find that the audits are looser for larger datasets and larger gradient clipping norms. We believe this is due to more noise reducing the empirical privacy leakage. When considering smaller datasets with 1,000 samples, our audits are tighter: for $\varepsilon=10.0$, we obtain $\varepsilon_{emp}=7.21$ and $6.95$, respectively, for MNIST and CIFAR-10.
@@
Overall, the ability to perform rigorous audits of differentially private machine learning techniques provides critical diagnostic tools to verify the correctness of implementations and libraries; moreover, it enables us to shed light on the tightness of theoretical guarantees in different settings.
@@ 2 Related Work
Membership Inference Attacks (MIAs). Shokri et al. [30] propose the first black-box MIAs against machine learning models using shadow models. Carlini et al. [7] improve by fitting a Gaussian distribution on the loss of the target sample and using fewer models to achieve equivalent attack accuracies. Sablayrolles et al. [29] consider per-example thresholds to improve the performance of the attacks in a low false-positive rate setting [7]. Finally, by factoring in various uncertainties in the MIA “game,” Ye et al. [35] construct an optimal generic attack. As mentioned, our work audits DP-SGD using state-of-the-art MIAs to estimate the empirical worst-case privacy leakage.
Auditing DP-SGD. Jayaraman and Evans [19] derive empirical privacy guarantees from DP-SGD, observing a large gap from the theoretical upper bounds, Jagielski et al. [18] use data poisoning in the black-box model to audit Logistic Regression and Fully Connected Neural Network models and produce tighter audits than using MIAs. They initialize models to fixed (average-case) initial parameters, but the resulting empirical privacy guarantees are still far from the theoretical bounds. They also adapt their procedure to the fine-tuning regime, again producing loose empirical estimates.
@@
In our work, we go beyond prior work by considering worst-case initial model parameters, which ultimately enables us to achieve significantly tighter audits of DP-SGD in the black-box threat model.
@@ 3.2 Differentially Private Machine Learning (DPML)
DP-SGD. Differentially Private Stochastic Gradient Descent (DP-SGD) [1] is a popular algorithm for training machine learning models while satisfying DP. In this paper, we consider classification tasks on samples from the domain $\mathcal{X}\times\mathcal{Y}$, where $\mathcal{X}$ is the features domain and $\mathcal{Y}$ is the labels domain.
DP-SGD, reviewed in Algorithm 1, introduces two hyper-parameters: gradient clipping norm $C$ and noise multiplier $\sigma$. Typically, $C$ is set to 1, while $\sigma$ is calibrated to the required privacy level $(\varepsilon,\delta)$ based on the number of iterations, $T$, and batch size, $B$. Early DP-SGD work [1] set batch sizes to be small and similar to those of non-private SGD; however, recent work has highlighted the advantage of setting large batch sizes [10, 13] by training state-of-the-art models with them. Following this trend, and to ease auditing, we set $B=N$, as also done in prior work [26]. Finally, while the initial set of model parameters $\theta_{0}$ is typically sampled randomly (e.g., using Xavier initialization [16]), it can also be set arbitrarily, without affecting the privacy guarantees provided by DP-SGD.
Algorithm 1 Differentially Private Stochastic Gradient
Descent (DP-SGD) [1]
@@
Fine-tuning regime. In this setting, one first pre-trains a model on a (large) public dataset and then fine-tunes it on the private data. When a suitable public dataset is available, prior work has shown that fine-tuning can achieve much higher utility than training from scratch using DP-SGD [10].
@@ 3.3 DP Auditing
Deriving $\varepsilon_{emp}$. When auditing a DP mechanism $\mathcal{M}$, one runs $\mathcal{M}$ repeatedly on two neighboring inputs $D$ and $D^{\prime}$ and the outputs $\left(O=\{o_{1},...,o_{R}\}\text{ and }O^{\prime}=\{o^{\prime}_{1},...,o^{\prime}_{R}\}\right)$ are given to an adversary $\mathcal{A}$. Next, $\mathcal{A}$ tries to distinguish between $O$ and $O^{\prime}$, which determines a false positive rate (FPR) and false negative rate (FNR). For a given number of outputs $2R$ and confidence level $\alpha$, empirical upper bounds $\overline{\text{FPR}}$ and $\overline{\text{FNR}}$ can be computed using Clopper-Pearson confidence intervals [27, 26]. Finally, the empirical upper bounds are converted to an empirical lower bound $\varepsilon_{emp}$ using the privacy region of $\mathcal{M}$ (see below) and $\varepsilon_{emp}$ can be compared to the theoretical $\varepsilon$. We consider an audit tight if $\varepsilon_{emp}\approx\varepsilon$. To ease presentation, we abstract away the details and use $\varepsilon_{emp}\leftarrow\text{EstimateEps}(\text{FPR},\text{FNR},2R,\alpha,\delta)$ to denote the process of estimating an $\varepsilon_{emp}$ from a given FPR and FNR.
Privacy region of DP-SGD. Given a mechanism $\mathcal{M}$, the privacy region of the mechanism, $\mathcal{R}_{\mathcal{M}}(\varepsilon,\delta)=\{(\text{FPR},\text{FNR})|\cdot\}$ defines the FPR and FNR values attainable for any adversary aiming to distinguish between $\mathcal{M}(D)$ and $\mathcal{M}(D^{\prime})$. Nasr et al. [26] note that the privacy region of DP-SGD corresponds (tightly) to that of $\mu$-Gaussian Differential Privacy [12], which can be used to first compute a lower bound on $\mu$:
@@ 4 Auditing Procedure
Threat model. We consider a simple black-box model where the adversary only sees the final model parameters, $\theta_{T}$, from DP-SGD. Specifically, unlike prior work on tight audits [26, 27], the adversary cannot observe intermediate model parameters or insert canary gradients. We do so to consider more realistic adversaries that may exist in practice. Additionally, we assume that the adversary can choose a worst-case target sample as is standard for auditing DP mechanisms [18].
Auditing. Our DP-SGD auditing procedure is outlined in Algorithm 2. As we operate in the black-box threat model, the procedure only makes a call to the DP-SGD algorithm—once again, they cannot insert canary gradients, etc. As mentioned, the adversary can set the initial model parameters ($\theta_{0}$), which is fixed for all models. Finally, we assume that the final model parameters are given to the adversary, who calculates the loss of the target sample as “observations.” These are then thresholded to compute FPRs and FNRs and estimate the $\varepsilon_{emp}$ using the EstimateEps process (see Section 3.3). This threshold, which we denote as $\tau$, must be computed on a separate set of observations (e.g., a validation set) for the $\varepsilon_{emp}$ to constitute a technically valid lower bound. However, as any decision threshold for GDP is equally likely to maximize $\varepsilon_{emp}$ [26], we follow common practice [24, 27, 26] to find the threshold that maximizes the value of $\varepsilon_{emp}$ from $O\cup O^{\prime}$.
@@
Crafting worst-case initial parameters. Recall that DP-SGD’s privacy analysis holds not only for randomly initialized models but also for arbitrary fixed parameters. Indeed, this was noted in prior work [10, 18, 26] and led to using initial model parameters that are fixed for all models. In our work, we consider worst-case, rather than average-case, initial parameters that minimize the gradients of normal samples in the dataset. This makes the target sample’s gradient much more distinguishable. To do so, we pre-train the model on an auxiliary dataset that follows the same distribution as the target dataset. Ultimately, we show that the gradients of normal samples are indeed minimized, achieving significantly tighter audits of DP-SGD in the black-box model.
@@ 5 Experiments
Due to computational constraints, we only train $2002R=$ 200 models (100 with the target sample and 100 without) to determine the empirical privacy estimates $\varepsilon_{emp}$ in each setting. In theory, we could achieve tighter audits by increasing the number of models, however, this would require significantly more computational resources and time (see below). Since our work focuses on the (broader) research questions of whether tight audits are possible in the black-box setting and how various factors affect this tightness, we believe $2002R=$ 200 provides a reasonable tradeoff between tightness and computational overhead.
In our evaluation, we report lower bounds with 95% confidence (Clopper-Pearson [9]) and report the mean and standard deviation values of $\varepsilon_{emp}$ over five independent runs.

### v3

@@ Abstract
This paper presents an auditing procedure for the Differentially Private Stochastic Gradient Descent (DP-SGD) algorithm in the black-box threat model that is substantially tighter than prior work. The main intuition is to craft worst-case initial model parameters, as DP-SGD’s privacy analysis is agnostic to the choice of the initial model parameters. For models trained on MNIST and CIFAR-10 at theoretical $\varepsilon=10.0$, our auditing procedure yields empirical estimates of $\varepsilon_{emp}=7.21$ and $6.95$, respectively, on a 1,000-record sample and $\varepsilon_{emp}=6.48$ and $4.96$ on the full datasets. By contrast, previous audits were only (relatively) tight in stronger white-box models, where the adversary can access the model’s inner parameters and insert arbitrary gradients. Overall, our auditing procedure can offer valuable insight into how the privacy analysis of DP-SGD could be improved and detect bugs and DP violations in real-world implementations. The source code needed to reproduce our experiments is available here.
@@ 1 Introduction
Differentially Private Stochastic Gradient Descent (DP-SGD) [1] allows training machine learning models while providing formal Differential Privacy (DP) guarantees [17]. DP-SGD is increasingly deployed in production systems [2, 5, 30] and supported in open-source libraries like Opacus [43], TensorFlow [21], and JAX [6]. Progress in privacy accounting and careful hyper-parameter tuning has helped reduce the gap in model performance compared to non-private training [13]. Furthermore, pre-training models on large public datasets and then fine-tuning them on private datasets using DP-SGD can also be used to improve utility [13].
@@
Implementing DP-SGD correctly can be challenging – as with most DP algorithms [7, 28] – and bugs that lead to DP violations have been found in several DP-SGD implementations [9, 24]. These violations do not only break the formal DP guarantees but can also result in realistic privacy threats [19, 36]. This prompts the need to audit DP-SGD implementations to verify whether the theoretical DP guarantees hold in practice. This usually involves running membership inference attacks [35], whereby an adversary infers whether or not a sample was used to train the model. The adversary’s success estimates the empirical privacy leakage from DP-SGD, and that is compared to the theoretical DP bound. If the former is “far” from the latter, it means that the auditing procedure is loose and may not be exploiting the maximum possible privacy leakage. Whereas, when empirical and theoretical estimates are “close,” auditing is tight and can be used to identify bugs and violations [31].
@@
Although techniques to tightly audit DP-SGD exist in literature [31, 32], they do so only in active white-box threat models, where the adversary can observe and insert arbitrary gradients into the intermediate DP-SGD steps. By contrast, in the more restrictive black-box threat model, where an adversary can only insert an input canary and observe the final trained model, prior work has only achieved loose audits [22, 31, 32]. In fact, prior work has suggested that the privacy analysis of DP-SGD in the black-box threat model can be tightened but only in limited settings (i.e., strongly convex smooth loss functions) [4, 11, 42]. Therefore, it has so far remained an open research question [31] whether or not it is possible to tightly audit DP-SGD in the black-box model.
@@
In this paper, we introduce a novel auditing procedure that achieves substantially tighter black-box audits of DP-SGD than prior work, even for natural (i.e., not adversarially crafted) datasets. The main intuition behind our auditing procedure is to craft worst-case initial model parameters, as DP-SGD’s privacy analysis is agnostic to the choice of initial model parameters. We show that the gap between worst-case empirical and theoretical privacy leakage in the black-box threat model is much smaller than previously thought [22, 31, 32]. When auditing a convolutional neural network trained on MNIST and CIFAR-10 at theoretical $\varepsilon=10$, our audits achieve an empirical privacy leakage estimate of $\varepsilon_{emp}=6.48$ and $4.96$, respectively, compared to $\varepsilon_{emp}=3.41$ and $0.69$ when the models were initialized to average-case initial model parameters, as done in prior work [13, 22, 31].
@@
In the process, we identify and analyze the key factors affecting the tightness of black-box auditing: 1) dataset size and 2) gradient clipping norm. Specifically, we find that the audits are looser for larger datasets and larger gradient clipping norms. We believe this is due to more noise, which reduces the empirical privacy leakage. When considering smaller datasets with 1,000 samples, our audits are tighter – for $\varepsilon=10.0$, we obtain $\varepsilon_{emp}=7.21$ and $6.95$, respectively, for MNIST and CIFAR-10.
@@
Overall, the ability to perform rigorous audits of differentially private machine learning techniques allows shedding light on the tightness of theoretical guarantees in different settings. At the same time, it provides critical diagnostic tools to verify the correctness of implementations and libraries.
@@ 2 Related Work
Membership Inference Attacks (MIAs).
Shokri et al. [35] propose the first black-box MIAs against machine learning models using shadow models. Carlini et al. [8] improve by fitting a Gaussian distribution on the loss of the target sample and using fewer models to achieve equivalent attack accuracies. Sablayrolles et al. [34] consider per-example thresholds to improve the performance of the attacks in a low false-positive rate setting [8]. By factoring in various uncertainties in the MIA “game,” Ye et al. [41] construct an optimal generic attack. Concurrent work [18, 27, 40] use poisoned pre-trained models, which is somewhat similar to our approach of using worst-case initial model parameters, but with a different focus. Specifically, Liu et al. [27] and Wen et al. [40] show that privacy leakage from MIAs is amplified when pre-trained models are poisoned but, crucially, do not consider DP auditing. Additionally, while Feng and Tramèr [18] achieve tight audits with poisoned pre-trained models, their strategy is geared towards the last-layer only fine-tuning regime. By contrast, we focus on auditing DP-SGD with full model training to estimate the empirical worst-case privacy leakage in the black-box setting.
Auditing DP-SGD.
Jayaraman and Evans [23] derive empirical privacy guarantees from DP-SGD, observing a large gap from the theoretical upper bounds. Jagielski et al. [22] use data poisoning in the black-box model to audit Logistic Regression and Fully Connected Neural Network models and produce tighter audits than using MIAs. They initialize models to fixed (average-case) initial parameters, but the resulting empirical privacy guarantees are still far from the theoretical bounds. They also adapt their procedure to the fine-tuning regime, again producing loose empirical estimates.
@@
Overall, while prior work [13, 22, 32] showed that reducing the randomness in the initial model parameters results in tighter empirical privacy leakage estimates, the audits remained loose as the fixed model parameters were still initialized randomly. Specifically, designing an effective adversarial strategy that provides tight audits in the black-box setting without destroying utility has remained, to the best of our knowledge, an open problem.
@@ 3.2 Differentially Private Machine Learning (DPML)
DP-SGD.
Differentially Private Stochastic Gradient Descent (DP-SGD) [1] is a popular algorithm for training machine learning models while satisfying DP. In this paper, we consider classification tasks on samples from the domain $\mathcal{X}\times\mathcal{Y}$, where $\mathcal{X}$ is the features domain and $\mathcal{Y}$ is the labels domain.
DP-SGD, reviewed in Algorithm 1, introduces two hyper-parameters: gradient clipping norm $C$ and noise multiplier $\sigma$. Typically, $C$ is set to 1, while $\sigma$ is calibrated to the required privacy level $(\varepsilon,\delta)$ based on the number of iterations, $T$, and batch size, $B$. Early DP-SGD work [1] set batch sizes to be small and similar to those of non-private SGD; however, recent work has highlighted the advantage of setting large batch sizes [13, 16] by training state-of-the-art models with them. Following this trend, and to ease auditing, we set $B=N$, as also done in prior work [31]. Finally, while the initial set of model parameters $\theta_{0}$ is typically sampled randomly (e.g., using Xavier initialization [20]), it can also be set arbitrarily (and even adversarially), without affecting the privacy guarantees provided by DP-SGD.
Algorithm 1 Differentially Private Stochastic Gradient Descent (DP-SGD) [1]
@@
Fine-tuning regime.
In this setting, one first pre-trains a model on a (large) public dataset and then fine-tunes it on the private data. When a suitable public dataset is available, prior work has shown that fine-tuning can achieve much higher utility than training from scratch using DP-SGD [13].
@@ 3.3 DP Auditing
Deriving $\varepsilon_{emp}$.
When auditing a DP mechanism $\mathcal{M}$, one runs $\mathcal{M}$ repeatedly on two neighboring inputs, $D$ and $D^{\prime}$, $R$ times each and the outputs $\left(O=\{o_{1},...,o_{R}\}\text{ and }O^{\prime}=\{o^{\prime}_{1},...,o^{\prime}_{R}\}\right)$ are given to an adversary $\mathcal{A}$. Next, $\mathcal{A}$ tries to distinguish between $O$ and $O^{\prime}$, which determines a false positive rate (FPR) and false negative rate (FNR). For a given number of outputs $2R$ and confidence level $\alpha$, empirical upper bounds $\overline{\text{FPR}}$ and $\overline{\text{FNR}}$ can be computed using Clopper-Pearson confidence intervals [32, 31]. Finally, the empirical upper bounds are converted to an empirical lower bound $\varepsilon_{emp}$ using the privacy region of $\mathcal{M}$ (see below) and $\varepsilon_{emp}$ can be compared to the theoretical $\varepsilon$. We consider an audit tight if $\varepsilon_{emp}\approx\varepsilon$. To ease presentation, we abstract away the details and use $\varepsilon_{emp}\leftarrow\text{EstimateEps}(\text{FPR},\text{FNR},2R,\alpha,\delta)$ to denote the process of estimating an $\varepsilon_{emp}$ from a given FPR and FNR. [^1] Please refer to https://github.com/spalabucr/bb-audit-dpsgd for details.
Privacy region of DP-SGD.
Given a mechanism $\mathcal{M}$, the privacy region of the mechanism, $\mathcal{R}_{\mathcal{M}}(\varepsilon,\delta)=\{(\text{FPR},\text{FNR})|\cdot\}$ defines the FPR and FNR values attainable for any adversary aiming to distinguish between $\mathcal{M}(D)$ and $\mathcal{M}(D^{\prime})$. Nasr et al. [31] note that the privacy region of DP-SGD corresponds (tightly) to that of $\mu$-Gaussian Differential Privacy [15], which can be used to first compute a lower bound on $\mu$:
@@ 4 Auditing Procedure
Threat model.
We consider a simple black-box model where the adversary only sees the final model parameters, $\theta_{T}$, from DP-SGD. Specifically, unlike prior work on tight audits [31, 32], the adversary cannot observe intermediate model parameters or insert canary gradients. We do so to consider more realistic adversaries that may exist in practice. Additionally, we assume that the adversary can choose a worst-case target sample as is standard for auditing DP mechanisms [22]. Finally, we assume that the adversary can choose (possibly adversarial) initial model parameters. Note that this is not only an assumption allowed by DP-SGD’s privacy analysis, but, in the context of the popular public pre-training, private fine-tuning regime [13], this is also a reasonable assumption.
Auditing.
Our DP-SGD auditing procedure is outlined in Algorithm 2. As we operate in the black-box threat model, the procedure only makes a call to the DP-SGD algorithm—once again, they cannot insert canary gradients, etc. As mentioned, the adversary can set the initial model parameters ($\theta_{0}$), which is fixed for all models. Finally, we assume that the final model parameters are given to the adversary, who calculates the loss of the target sample as “observations.” These are then thresholded to compute FPRs and FNRs and estimate the $\varepsilon_{emp}$ using the EstimateEps process (see Section 3.3). This threshold, which we denote as $\tau$, must be computed on a separate set of observations (e.g., a validation set) for the $\varepsilon_{emp}$ to constitute a technically valid lower bound. However, as any decision threshold for GDP is equally likely to maximize $\varepsilon_{emp}$ [31], we follow common practice [29, 32, 31] to find the threshold that maximizes the value of $\varepsilon_{emp}$ from $O\cup O^{\prime}$.
@@
Crafting worst-case initial parameters.
Recall that DP-SGD’s privacy analysis holds not only for randomly initialized models but also for arbitrary fixed parameters. Indeed, this was noted in prior work [13, 22, 31] and led to using initial model parameters that are fixed for all models. In our work, we consider worst-case, rather than average-case, initial parameters that minimize the gradients of normal samples in the dataset. This makes the target sample’s gradient much more distinguishable. To do so, we pre-train the model using non-private SGD on an auxiliary dataset that follows the same distribution as the target dataset. Specifically, for MNIST, we pre-train the model on half of the full dataset (reserving the other half to be used for DP-SGD) for 5 epochs with a batch size of 32 and a learning rate of 0.01. For CIFAR-10, we first pre-train the model on the CIFAR-100 dataset for 300 epochs with batch size 128 and a learning rate of 0.1. Then, we (non-privately) fine-tune the model on half of the full dataset for 100 epochs, with a batch size of 256 and a learning rate of 0.1. Ultimately, we show that the gradients of normal samples are indeed minimized, achieving significantly tighter audits of DP-SGD in the black-box model.
@@ 5 Experiments
Due to computational constraints, we only train $2002R=$ 200 models (100 with the target sample and 100 without) to determine the empirical privacy estimates $\varepsilon_{emp}$ in the full model training setting and train $10002R=$ 1000 models for the last-layer only fine-tuning setting. In theory, we could achieve tighter audits by increasing the number of models, however, this would require significantly more computational resources and time (see below). Since our work focuses on the (broader) research questions of whether tight audits are possible in the black-box setting and how various factors affect this tightness, we believe $2002R=$ 200 provides a reasonable tradeoff between tightness and computational overhead.
In our evaluation, we report lower bounds with 95% confidence (Clopper-Pearson [12]) and report the mean and standard deviation values of $\varepsilon_{emp}$ over five independent runs. [^2] The source code needed to reproduce our experiments is available from https://github.com/spalabucr/bb-audit-dpsgd.